OpenStudy (anonymous):

Calculate the derivatives of the following real functions by using the chain rule; maybe several times. In each case, determine the inner and outer function to be used, together with their respective areas of definition. d) \(h(x) = \sqrt{2+exp(x^{2})}\)

OpenStudy (anonymous):

derivative of outside function, evaluated at the inside function, times the derivative of the inside function

OpenStudy (anonymous):

hey @satellite73 thank you but i am not so good at math, pls help me more, in about 5 hours i need to give up homework :) and i can write you a message unfortunately..its blocked i think.

OpenStudy (anonymous):

*cant

OpenStudy (turingtest):

I refuse to accept the label "not good at math" especially when someone has made it to calculus

OpenStudy (anonymous):

the derivative of the square root of something is one of two times root of something, times the derivative of something

OpenStudy (anonymous):

more succinctly written as \[\frac{d}{dx}\sqrt{f(x)}=\frac{f'(x)}{2\sqrt{f(x)}}\]

OpenStudy (anonymous):

@TuringTest thank you, i didnt really understand what you mean but i think you are saying something good (my english is bad as my math)

OpenStudy (turingtest):

no worries, I'm just saying that you should not decide that you are bad at math so quickly

OpenStudy (anonymous):

aah ok, i understood like that but i was not sure :) thanks, i need to be good at math very soon i have exam next month, and i had always problem wit this lecture in germany we are sayng its a analysis i think in america its calculus

OpenStudy (anonymous):

so here your \(f(x)=2+\exp(x^2)\) and so your \(f'(x)=2x\exp(x^2)\)

OpenStudy (turingtest):

math analysis=calculus in English too :)

OpenStudy (anonymous):

and therefore you answer is \[\frac{2x\exp(x^2)}{2\sqrt{2+\exp(x^2)}}\] make sure to cancel the 2 so your professor does not think you are daft

OpenStudy (anonymous):

@satellite73 ok.. which 2 you mean? @TuringTest but i think this time i can do it with exam becuase i found openstudy here so many people who like to help, but the problem is until this days i didnt had much time to learn,

OpenStudy (zarkon):

\[\frac{\cancel{2}x\exp(x^2)}{\cancel{2}\sqrt{2+\exp(x^2)}}\]

OpenStudy (anonymous):

aah ok thanks Zarkon

OpenStudy (turingtest):

Please keep coming to Open Study for help, that is our goal @satellite73 and @Zarkon are some of the best, so they can sure help you :)

OpenStudy (anonymous):

i love \(\cancel{cancel}\)

OpenStudy (zarkon):

I'm lazy and don't help much Satellite is good though ;)

OpenStudy (anonymous):

@satellite73 thank you very much @TuringTest yes i know, Sattelite helping me much, Zarkon he helped first today but its ok :) sorry if i forget that you helped maybe

OpenStudy (turingtest):

Zarkon only provides the most minimal detail possible, which I love because it forces you to think. Satellite is more explanatory; just a different style I guess

OpenStudy (anonymous):

yes Zarkons help i will need before exam, when i learned and need details :)

OpenStudy (anonymous):

as you said

OpenStudy (zarkon):

I use the moore method...you do more and I do less ;)

OpenStudy (anonymous):

:D

OpenStudy (turingtest):

@Zarkon you have "giant footsteps" though, as asnaseer says. One hint from Zarkon is worth twenty from me.... in my opinion at least

OpenStudy (zarkon):

TT you are an excellent tutor...more often more helpful than I am.

OpenStudy (anonymous):

@TuringTest are you computer scientist or mathematician?

OpenStudy (turingtest):

praise from Caesar :D

OpenStudy (turingtest):

I am a lowly student

OpenStudy (zarkon):

for those that don't know there is a Moore method... http://en.wikipedia.org/wiki/Moore_method

OpenStudy (turingtest):

ha, I thought you just made that up!

OpenStudy (zarkon):

I took two classes where the Moore method was used...topology and geometry

OpenStudy (turingtest):

geometry? some super special Reimannian Manifold geometry or something o-0

OpenStudy (zarkon):

we built up euclidean geometry from the ground up. It was a senior level undergraduate class.

OpenStudy (zarkon):

the prof just gave definitions and axioms...and the students had to prove all the theorems

OpenStudy (zarkon):

no lecture

OpenStudy (anonymous):

are you both in same university ?

OpenStudy (turingtest):

cool :) that's I had a similar experience in geometry in high school that led me to feel confident in math, but that story is for another time :) and no @tunahan , Zarkon is a master, a real pro I'm just a student

OpenStudy (zarkon):

I am no longer a student.

OpenStudy (turingtest):

...and we live in different countries, etc. I just like to tutor

OpenStudy (anonymous):

ok :) i thought the most of the people here is from usa

OpenStudy (anonymous):

i thought even they are the most from same university..

OpenStudy (turingtest):

we both are, but I live in mexico but this site has a nice mix of US, European, middle-eastern and asian students

OpenStudy (anonymous):

ok this site is really helpfull, if see who made it, i would like to thank him or people who has done..

OpenStudy (zarkon):

Some math professors also stop by to help

OpenStudy (turingtest):

@tunahan I'm glad you find value in the purpose of this site. I hope you continue to come here, and become an integral part of OS you could really contribute with that attitude!

OpenStudy (anonymous):

yes this site is really like a treausre for me, OS is not Operating System right :) ?

OpenStudy (turingtest):

Not in this case OS=Open Study (sorry, I just type it so many times a day...) anyway, I've got to go pleasure talking to you, see you around!

OpenStudy (anonymous):

sorry i am stupid :) i would know that u mean Open Study, i am computer science student, and mad with operating systems, ok it was pleausre to talking with you Turing, i have more questions today, you can both take a look if you would like to

OpenStudy (anonymous):

@Zarkon here is my question may be you would like to solve for me so that i get good points from my mentor :) http://openstudy.com/study#/updates/4fe3d9aae4b06e92b8729084